OpenStudy (anonymous):

Kenji has $600 to spend on exercise classes. Each weight training class(x) costs $20 and each aerobic class(y) costs $15. Write an inequality for this situation and determine which ordered pair is a solution to the inequality

OpenStudy (texaschic101):

20x + 15y <= 600 one solution is : (15,20)

OpenStudy (anonymous):

thats not an answer choice

OpenStudy (texaschic101):

give me your answer choices

OpenStudy (anonymous):

25,20 10,35 10,20 30,20

OpenStudy (texaschic101):

20x + 15y <= 600 (25,20) 20(25) + 15(20) <= 600 500 + 300 <= 600 800 <= 600 (not this one) (10,35) 20x + 15y <= 600 20(10) +15(35) <= 600 200 + 525 <= 600 725 <= 600 (not this one) (10,20) 20x + 15y <= 600 20(10) + 15(20) <= 600 200 + 300 <= 600 500 <= 600 (this one works) (30,20) 20x + 15y <= 600 20(30) + 15(20) <= 600 600 + 300 <= 600 900 <= 600 (not gonna work answer is : (10,20)